the area of a rectangular field is 2275 . The field is enclosed by 200 m of fencing. What are the dimensions of the field
Answer by flame8855(424)

xy=2275
x+y = 100
2275/x+x=100
2275+x^2 -100x =0
x= 65
x = 32
if x = 65 y = 35
if x = 32 y = 68